OpenStudy (anonymous):

I'm really confused!!! Can someone PLEASE help me?!?! Solve the literal equation P=2L+2W for the unknown variable. Be sure to show all the steps. P=920 L=100

OpenStudy (anonymous):

p = 2l + 2w 2w = p - 2l w = p-2L divided by 2

OpenStudy (anonymous):

plug L in, and then plug in P. multiply L x 2. then subtract on both sides. after that you get 720=2W. then divide by two and it should equal 360 answer:360

OpenStudy (anonymous):

Thanks!